# TyphoonMLA: A Mixed Naive-Absorb MLA Kernel For Shared Prefix

### v1

@@ 4 Experiments
Figure 5: Latency breakdown of TyphoonMLA (bars on the left-hand side) and CATLASS absorb-only baseline (bars on the right-hand side). Stage 1 and Stage 2 represent the naive and absorb parts of TyphoonMLA, respectively.
HBM Footprint.
Since TyphoonMLA stores the shared portion of the KV-cache in an uncompressed form, its memory footprint differs from that of the absorb baselines. To quantify the impact of TyphoonMLA on the memory footprint, we analyze TyphoonMLA’s HBM usage under various deployment settings and compare it against the absorb baseline. We assume that the model is distributed across 384 NPUs, as in a CloudMatrix cluster (Zuo et al., 2025), employing full expert parallelism on MoE layers and a combination of data, tensor, and sequence parallelism of factors 24, 4, and 4, respectively. Fig.6 shows the HBM size of DeepSeek-v3 for batch sizes ranging from 4K to 32K and maximum sequence lengths from 32K to 256K, assuming Prompt A as the shared prefix.
@@ 5 Related Work
LLM Serving Acceleration. Various prior work proposes system-level optimizations for LLM serving. Orca (Yu et al., 2022) introduced continuous batching, which allows replacing the completed requests in a batch with new ones to improve effective throughput. Inspired by the OS virtual memory concept, PagedAttention (Kwon et al., 2023) partitions the KV-cache into memory pages to efficiently handle KV-cache with variable length. TyphoonMLA supports both continuous batching and PagedAttention; thus, it benefits from the performance gains of such system-level optimizations. Furthermore, to distribute LLM inference across multiple devices, existing systems typically employ tensor (Shoeybi et al., 2019) and sequence (Jacobs et al., 2023) parallelism. TyphoonMLA also seamlessly supports both types of parallelization, enabling deployment at scale.
Efficient Attention Kernels. As attention computation constitutes a significant portion of the total execution time of LLM inference, many prior studies focused on developing efficient attention kernels. Flash Attention (Dao et al., 2022) partitions attention computations into tiles that fit on the on-chip buffers to minimize HBM bandwidth usage. FlashDecoding (Dao, 2024) introduces the Split-K method to improve the parallelization across GPU cores. Flashinfer (Ye et al., 2025) provides a customizable attention template to facilitate the development of custom attention kernels. However, all these techniques are developed for MHA; hence, they are not applicable to MLA. FlashMLA (Jiashi Li, 2025), FlashMLA-ETAP (Dege et al., 2025), and ThunderMLA (Spector et al., 2025) propose efficient kernel designs for MLA. Unfortunately, these kernels are agnostic to the shared prefix and an absorb-only implementation; therefore, they can not fully exploit hardware resources.
Prefix Sharing. Several prior works have proposed techniques to exploit shared prefixes in attention calculations and KV-cache management (Juravsky et al., 2024; Zheng et al., 2024). Other studies developed specialized GPU kernels for tree-structured KV-caches to leverage the data reuse introduced by shared prefixes (Ye et al., 2024a; Zhu et al., 2024; Ye et al., 2024b; Yao et al., 2025; Pan et al., 2025; Wang et al., 2025b). However, these kernel designs target MHA and aim to reduce HBM bandwidth usage. Consequently, they are not applicable to MLA, whose decode stage is typically compute-bound. In contrast, TyphoonMLA addresses this gap by reducing the number of FLOPS required in the compute-bound decode stage of MLA, achieving substantial speedups.
@@ 6 Conclusions
In this paper, we proposed a novel MLA method, TyphoonMLA, which combines the naive and absorb implementations to effectively exploit shared prefixes in the KV-cache. Our theoretical analysis showed that, in the presence of a shared prefix, TyphoonMLA requires fewer FLOPS in compute-bound parts and smaller HBM bandwidth in memory-bound parts of the attention calculations than the existing MLA kernels. Experimental results on NPUs and GPUs show that TyphoonMLA improves attention throughput by up to 3.2× on DeepSeek-v3 and Kimi K2 models, while incurring only a minimal increase in HBM footprint. These results confirm that TyphoonMLA offers a practical and effective solution to improve the performance and efficiency of LLM inference, enhancing the user experience and reducing the cost of LLM applications.
@@ Appendix A Appendix
A.1 Theoretical Analysis of TyphoonMLA
@@
A.2 Sensitivity to batch size

### v2

@@ 4 Experiments
Figure 4: Latency breakdown of TyphoonMLA (bars on the left-hand side) and CATLASS absorb-only baseline (bars on the right-hand side) for Kimi K2 architecture. Stage 1 and Stage 2 represent the naive and absorb parts of TyphoonMLA, and have sequence lengths of 4096 and 512, respectively.
End-to-End Speedup: Previous experimental results demonstrated that TyphoonMLA significantly accelerates the attention layers in MLA architectures. However, the execution time of the other layers in LLMs (e.g., MoE, collective communication etc.) would remain unchanged. Therefore, we now investigate the impact of the proposed method on the end-to-end LLM decoding performance. To that end, we estimate the token generation rate (TGR) for DeepSeek-v3 with a batch size of 128 per device on a system with 128 GPUs by measuring attention layers on a GPU and using the profiling data for other layers provided by DeepSeek-AI (Deepseek-AI, 2025).
Table 3 reports the estimated TGR and per-iteration execution time for both the attention layers and the full model when using TyphoonMLA, comparing against those with FlashMLA as the baseline. Since attention computation constitutes a substantial portion of overall runtime, the speedup achieved by TyphoonMLA in the attention layers translates directly into notable end-to-end throughput gains, reaching up to a 1.48× improvement in tokens per second.
HBM Footprint: Since TyphoonMLA stores the shared portion of the KV-cache in an uncompressed form, its memory footprint differs from that of the absorb baselines. To quantify the impact of TyphoonMLA on the memory footprint, we analyze TyphoonMLA’s HBM usage under various deployment settings and compare it against the absorb baseline. We assume that the model is distributed across 384 NPUs, as in a CloudMatrix cluster (Zuo et al., 2025), employing full expert parallelism on MoE layers and a combination of data, tensor, and sequence parallelism of factors 24, 4, and 4, respectively. Fig. 5 shows the HBM size of DeepSeek-v3 for batch sizes ranging from 4K to 32K and maximum sequence lengths from 32K to 256K, assuming Prompt A as the shared prefix.
@@ 5 Related Work
LLM Serving Acceleration: Various prior work proposes system-level optimizations for LLM serving. Orca (Yu et al., 2022) introduced continuous batching, which allows replacing the completed requests in a batch with new ones to improve effective throughput. Inspired by the OS virtual memory concept, PagedAttention (Kwon et al., 2023) partitions the KV-cache into memory pages to efficiently handle KV-cache with variable length. TyphoonMLA supports both continuous batching and PagedAttention; thus, it benefits from the performance gains of such system-level optimizations. Furthermore, to distribute LLM inference across multiple devices, existing systems typically employ tensor (Shoeybi et al., 2019) and sequence (Jacobs et al., 2023) parallelism. TyphoonMLA also seamlessly supports both types of parallelization, enabling deployment at scale.
Efficient Attention Kernels: As attention computation constitutes a significant portion of the total execution time of LLM inference, many prior studies focused on developing efficient attention kernels. Flash Attention (Dao et al., 2022) partitions attention computations into tiles that fit on the on-chip buffers to minimize HBM bandwidth usage. FlashDecoding (Dao, 2024) introduces the Split-K method to improve the parallelization across GPU cores. Flashinfer (Ye et al., 2025) provides a customizable attention template to facilitate the development of custom attention kernels. However, all these techniques are developed for MHA; hence, they are not applicable to MLA. FlashMLA (Jiashi Li, 2025), FlashMLA-ETAP (Dege et al., 2025), and ThunderMLA (Spector et al., 2025) propose efficient kernel designs for MLA. Unfortunately, these kernels are agnostic to the shared prefix and an absorb-only implementation; therefore, they can not fully exploit hardware resources.
Prefix Sharing: Several prior works (e.g., SGLang, Hydragen etc.) have proposed techniques to exploit shared prefixes in attention calculations and KV-cache management (Juravsky et al., 2024; Zheng et al., 2024). Other frameworks (e.g., FlashInfer-Cascade, FastTree etc.) developed specialized MHA/GQA GPU kernels for tree-structured KV-caches to leverage the data reuse introduced by shared prefixes (Ye et al., 2024a; Zhu et al., 2024; Ye et al., 2024b; Yao et al., 2025; Pan et al., 2025; Wang et al., 2025b). However, these kernel designs target MHA/GQA architectures and aim to reduce HBM bandwidth usage. Consequently, they are not applicable to MLA, whose decode stage is typically compute-bound. In contrast, TyphoonMLA addresses this gap by reducing the number of FLOPS required in the compute-bound decode stage of MLA, achieving substantial speedups. Moreover, the mechanisms to handle complex tree-structured KV-caches introduced by the prior work would be equally applicable to both naive and absorb implementations. Therefore, applying these techniques to TyphoonMLA would be trivial.
@@ 6 Conclusions
In this paper, we proposed a novel MLA method, TyphoonMLA, which combines the naive and absorb implementations to effectively exploit shared prefixes in the KV-cache. Our theoretical analysis showed that, in the presence of a shared prefix, TyphoonMLA requires fewer FLOPS in compute-bound parts and smaller HBM bandwidth in memory-bound parts of the attention calculations than the existing MLA kernels. Experimental results on NPUs and GPUs show that TyphoonMLA improves attention throughput by up to 3.2× on DeepSeek-v3 and Kimi K2 models, while incurring only a minimal increase in HBM footprint.
@@ Appendix A Appendix
A.1 Roofline Analysis
[…]
A.2 Theoretical Analysis of TyphoonMLA
@@
A.3 Sensitivity to batch size